A 12 kg mass was projected with

an initial kinetic energy of 450J along

the surface of a rough plane inclined

at 30° above the horizontal. If the

mass comes to rest after traveling 3.5m along the plane, calculate the

coefficient of kinetic friction between the mass and the plane.



1
Expert's answer
2021-01-31T03:40:10-0500

Let's apply the Newton's Second Law of Motion:


"-mgsin\theta-\mu_kmgcos\theta=ma."

From this equation we can find the coefficient of kinetic friction between the mass and the plane:


"\mu_k=\dfrac{-gsin\theta-a}{gcos\theta}."

Let's find the initial velocity of the mass from the definition of kinetic energy:


"v=\sqrt{\dfrac{2KE}{m}}=\sqrt{\dfrac{2\cdot450\ J}{12\ kg}}=8.66\ \dfrac{m}{s}."

Then, we can find the deceleration of the mass:


"a=\dfrac{v^2-v_0^2}{2d}=\dfrac{0-(8.66\ \dfrac{m}{s})^2}{2\cdot3.5\ m}=-10.71\ \dfrac{m}{s}."

Finally, substituting "a" into the formula for "\mu_k", we get:


"\mu_k=\dfrac{-9.8\ \dfrac{m}{s^2}\cdot sin30^{\circ}-(-10.71\ \dfrac{m}{s})}{9.8\ \dfrac{m}{s^2}\cdot cos30^{\circ}}=0.68"